OpenStudy (anonymous):

What is 8 1/5 - 4 2/5 equal please help me

OpenStudy (anonymous):

yu would want to convert 8 1/5 to a improper fraction and do the same for 4 2/5, then yu subtract them and wold result in finding the lcm

OpenStudy (anonymous):

the answer would be 19

OpenStudy (anonymous):

someone help me

OpenStudy (anonymous):

its like this 8*5+1/5=41/5 4*5+2/5=22/5 41/5-22/5 19/5

OpenStudy (anonymous):

so would put tht into a mixed number cause i can't leave it improper

OpenStudy (anonymous):

yes . . . its how its solved. . .

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

it would be 3 4/5 right